# Gagan Bhatia      El Moatez Billah Nagoudi     Hasan Cavusoglu Muhammad Abdul-Mageed

### v1

@@ Abstract
We introduce FinTral, a suite of state-of-the-art multimodal large language models (LLMs) built upon the Mistral-7b model and tailored for financial analysis. FinTral integrates textual, numerical, tabular, and image data. We enhance  FinTral with domain-specific pretraining, instruction fine-tuning, and RLAIF training by exploiting a large collection of textual and visual datasets we curate for this work. We also introduce an extensive benchmark featuring nine tasks and 25 datasets for evaluation, including hallucinations in the financial domain. Our  FinTral model trained with direct preference optimization employing advanced Tools and Retrieval methods, dubbed FinTral-DPO-T&R, demonstrates an exceptional zero-shot performance. It outperforms ChatGPT-3.5 in all tasks and surpasses GPT-4 in five out of nine tasks, marking a significant advancement in AI-driven financial technology. We also demonstrate that FinTral has the potential to excel in real-time analysis and decision-making in diverse financial contexts.
@@ 1 Introduction
The rest of this paper is organized as follows: In Section 2, we review related work with a particular emphasis on financial LLMs, their applications and challenges. Section 3 outlines how we built our benchmark dataset: FinSet. We present our approach to model pretraining, instruction tuning, and prompting strategies, and subsequently introduce FinTral models in Section 4. In Section 5, we present our experiments and comprehensively analyse our models. We discuss our results in Section 6 and conclude in Section 7.
@@ 3.1 Pretraining Dataset
We introduce FinSet, a 20 billion token, high quality dataset we build for financial LLM training. FinSet is acquired based on a collection of large text corpora (2.9 billion documents, making 135 billion tokens; see Table 1) from which we extract finance-specific data that we then clean using a careful filtering pipeline. The datasets are described in Appendix B. Our cleaning pipeline is detailed in Appendix C. Our document cutoff date is August 1, 2023, which affords recent information to our models.
@@ 4 Fintral
You are a financial expert specializing in the nuanced analysis of financial statements and a wide array of data-driven financial tasks. For each prompt you are given, think step by step. Sometimes, you must extract relevant information to proceed with the problem. Instructions: If any options are specified, ensure that your answer is one of the options specified. new Do not explain why you think the answer is correct.
@@ 5 Experiments
[…]
@@ 9 Ethics Statement
Data. Our pretraining datasets are collected from public domains, encompassing a wide range of financial topics and sources. While these datasets provide comprehensive coverage for financial language modeling, we must be aware of the potential biases and limitations inherent in publicly available data, ensuring our model remains as objective and unbiased as possible.
@@
Model Release. We plan to release our models responsibly. Given the sensitive nature of financial data and the potential for misuse, we will implement strict guidelines and conditions for the use of FinTral, particularly in real-world applications. This includes clear guidelines on ethical usage and the avoidance of deployment in contexts that could lead to unethical practices such as market manipulation or privacy violations.
@@ Appendix C Financial data cleaning and deduplication pipeline
We started of gathering various text corpora shown in table 1, resulting in a dataset consisting of 2.9B documents.
Figure C.1: Explanation of our data deduplication pipeline
The data that we collected is not only unclean but also suffers from large-scale duplication. As shown by Soboleva et al. (2023), using clean and deduplicated data is computationally efficient for model training. The data cleaning and deduplication pipeline for financial data begins with URL filtering, in which the raw data is initially processed. This crucial step ensures the inclusion of only pertinent URLs, enhancing the dataset’s quality by excluding irrelevant or unsuitable sources. Once the URLs are streamlined, the Text extraction phase commences, whereby contents of documents from the selected URLs are meticulously extracted, filtering out images while maintaining the large dataset scale. Following this, the language identification phase excluded non-English documents by categorizing them based on the language of their tokens. Subsequently, the pipeline further refines the data through Document-wise domain-based filtering, narrowing down to 100 billion tokens pertinent to the financial domain by excluding 55B-token non-financial documents. Recognizing the importance of data privacy and relevance, the pipeline incorporates removing sensitive information, which is done using a classifier built using FinBERT Araci (2019). Line-wise corrections enhance accuracy and filter out 5B tokens of sensitive information. An extensive Fuzzy deduplication process reduces the data to 38 billion tokens. This is followed by an Exact deduplication method, which trims another 13 billion tokens. Finally, the text cleaning process identifies and excludes 5B improper tokens, including all sensitive information. Ultimately, the pipeline crafts a streamlined financial dataset, culminating in a concise 20B-token financial dataset. The pipeline is illustrated in figure C.1.
@@ Appendix D Downstream Dataset details
Figure D.1: Illustration of the FinBEAT VL Model’s Proficiency in Chart Understanding Tasks. The panel demonstrates the model’s ability to interpret stock price movements and identify peak values from a candlestick chart.
@@
Retrieval Data extraction Answer Generation From the following PDF file, extract all the relevant information that might help in answering the question: PDF: URL Extracted Text: text…….. Question: ask question? You are a financial expert specializing in the nuanced analysis of financial statements and a wide array of data-driven financial tasks. And think step by step for each prompt. Sometimes, you must extract relevant information to proceed with the problem. Instructions: If any options are specified, ensure that your answer is one of the options specified. new Do not explain why you think the answer is correct.
new Answer the question by formulating your response using predefined mathematical functions. For addition, use Add(a, b), which represents a + b. Use Subtract(a, b) to denote a - b for subtraction. Construct your answer by combining these functions appropriately to reflect the required calculations. Context: text + table + image Retrieval: Retrieved relevant information Question: Task based question? Answer:
@@
You are a financial expert specializing in the nuanced analysis of financial statements and a wide array of data-driven financial tasks. And think step by step for each prompt. Sometimes, you must extract relevant information to proceed with the problem. Instructions: new If any options are specified, ensure that your answer is one of the options specified. new Do not explain why you think the answer is correct.
@@
Figure D.4: FinBEAT Prompting method Example
@@
Examples of No Hallucinations Prompt: What is the definition of ’Low/No Documentation Loan’ in the financial domain? Wikipedia: A No-Doc or Low-doc loan (abbr: No/Low Documentation Loan) refers to loans that do not require borrowers to provide income documentation to lenders or require little documentation. FinTral-RL-T&R: A low/no documentation loan allows a potential borrower to apply for a mortgage while providing little or no information regarding their employment, income, or assets. Regulation of these loans has evolved significantly since 2008, but they remain an option for some borrowers in nontraditional financial situations. Annotator Rating: A ChatGPT: A low/no documentation loan is a type of mortgage or loan where the borrower is not required to provide extensive financial documentation, such as income and asset verification, making it easier to qualify but often resulting in higher interest rates. Annotator Rating: A GPT-4: A Low/No Documentation Loan is a type of loan that requires minimal or no verification of the borrower’s income, assets, or employment, often used by individuals who may not have traditional income sources or documentation. Annotator Rating: A Examples of LLMs Hallucinations Prompt: What is the definition of ’Open Market’ in the financial domain? Wikipedia: An open market is an economic system without trade barriers to free market activities. In an open market, buyers and sellers can do business freely without common barriers, such as unfair licensing agreements, arbitrary taxes, unionization, subsidies, and other regulations affecting regular market operations. FinTral-RL-T&R: An open market is an economic system with little to no barriers to free-market activity. An open market is characterized by the absence of tariffs, taxes, licensing requirements, subsidies, unionization, and any other regulations or practices that interfere with free-market activity. Open markets may have competitive barriers to entry, but never any regulatory barriers to entry. Annotator Rating: A ChatGPT: In finance, the Open Market refers to a marketplace where securities are traded directly between buyers and sellers without the intervention of a third party, such as an exchange. Annotator Rating: C GPT-4: An "Open Market" in the financial domain refers to an economic system with few barriers to free market activity. It typically involves unrestricted buying and selling of goods and services, with prices determined by supply and demand, and minimal government intervention in terms of tariffs, quotas, or regulations. This concept is fundamental in the operation of capital markets, where securities are freely traded. Annotator Rating: A

### v2

@@ Abstract
We introduce FinTral, a suite of state-of-the-art multimodal large language models (LLMs) built upon the Mistral-7b model and tailored for financial analysis. FinTral integrates textual, numerical, tabular, and image data. We present FinSet, the largest financial LLM pretraining training, instruction tuning and financial alignment dataset and evaluation benchmark featuring nine tasks and 23 datasets and the first to understand and mitigate financial hallucinations. We enhance FinTral with domain-specific pretraining, instruction fine-tuning, and RLAIF training by exploiting a large collection of textual and visual datasets we curate for this work. We also introduce an extensive benchmark featuring nine tasks and 25 datasets for evaluation, including hallucinations in the financial domain. Our  FinTral model trained with direct preference optimization employing advanced Tools and Retrieval methods, dubbed FinTral-DPO-T&R, demonstrates an exceptional zero-shot performance. It outperforms ChatGPT-3.5 in all tasks and surpasses GPT-4 in five out of nine tasks, marking a significant advancement in AI-driven financial technology. We also demonstrate that FinTral has the potential to excel in real-time analysis and decision-making in diverse financial contexts. The GitHub repository for FinTral is available at https://github.com/UBC-NLP/fintral.
@@ 1 Introduction
The rest of this paper is organized as follows: In Section 2, we review related work with a particular emphasis on financial LLMs, their applications and challenges. Section 3 outlines how we built our benchmark dataset: FinSet. We present our approach to model pretraining, instruction tuning, and prompting strategies, and subsequently introduce FinTral models in Section 4. In Section 5, we present our experiments and comprehensively analyse our models. We discuss our results in Section 6 and conclude in Section 7. Finally, we discuss limitations in Section 8 and provide an ethics statement in Section 9.
@@ 3.1 Pretraining Dataset
Figure 3: Explanation of our data deduplication pipeline
We introduce FinSet, a 20 billion token, high quality dataset we build for financial LLM training. FinSet is acquired based on a collection of large text corpora (2.9 billion documents, making 135 billion tokens; see Table 1) from which we extract finance-specific data that we then clean using a careful filtering pipeline. The datasets are described in Appendix B. Our cleaning pipeline is detailed in Appendix C and is further illustrated in Figure 3. Our document cutoff date is August 1, 2023, which affords recent information to our models.
@@ 3.5 Downstream Evaluation Datasets
Figure 4: Illustration of the FinVQA Dataset to understand model’s proficiency in Chart Understanding Tasks. The figure demonstrates the model’s ability to interpret stock price movements and identify peak values from a candlestick chart.
@@ 4 Fintral
You are a financial expert specializing in the nuanced analysis of financial statements and a wide array of data-driven financial tasks. For each prompt you are given, think step by step. Sometimes, you must extract relevant information to proceed with the problem. Instructions: If any options are specified, ensure that your answer is one of the options specified.
new Do not explain why you think the answer is correct.
@@ 5 Experiments
We conducted multiple experiments to illustrate the efficacy of the methods described in section 4. We evaluated our model on the downstream tasks described in section 3.5. Symbols in the following tables indicate the types of models: $\clubsuit$, $\spadesuit$, $\diamondsuit$, $\heartsuit$, $\bigstar$ and, $\blacksquare$ represent the pre-trained model, the fine-tuned model, the instruction fine-tuned model, the RL-Tuned Models, tools, and retrieval, respectively. We then performed a hallucination index accuracy check to assess how well our model mitigates one of the biggest challenges for LLMs.
@@ 6 Discussion
6.1 Model Shortcomings Analysis
Each iteration of FinTral was designed to progressively address the shortcomings of its predecessors, focusing on enhancing financial domain-specific knowledge and instruction compliance while minimizing the occurrence of hallucinations.
The initial versions, starting from Mistral-7B and the subsequent FinTral, showed improved domain knowledge but struggled with maintaining instruction accuracy and often produced misleading outputs. FinTral-INST, despite advancements, continued to face challenges in adhering to specific task formats and response coherence. This was partly mitigated in FinTral-DPO, which better adhered to the instructed formats by introducing direct preference optimization.
The most advanced iteration, FinTral-DPO-T&R, integrates tools and retrieval capabilities to further refine performance, especially in complex financial tasks requiring extensive data integration and computational power. This version demonstrates substantial improvements in handling detailed financial analyses and proves highly competitive, even with leading models like GPT-4, particularly in domains requiring rigorous financial insight.
@@ 9 Ethics Statement
Data. Our pretraining datasets are collected from public domains, encompassing a wide range of financial topics and sources. While these datasets provide comprehensive coverage for financial language modelling, we must be aware of the potential biases and limitations inherent in publicly available data, ensuring our model remains as objective and unbiased as possible.
@@
Model Release. We are considering releasing our models and evaluation data (FinSET) responsibly. Given the sensitive nature of financial data and the potential for misuse, we will implement strict guidelines and conditions for the use of FinTral, particularly in real-world applications. This includes clear guidelines on ethical usage and the avoidance of deployment in contexts that could lead to unethical practices such as market manipulation or privacy violations.
@@ Appendix C Financial data cleaning and deduplication pipeline
We started of gathering various text corpora shown in table 1, resulting in a dataset consisting of 2.9B documents. The data that we collected is not only unclean but also suffers from large-scale duplication. As shown by Soboleva et al. (2023), using clean and deduplicated data is computationally efficient for model training. The data cleaning and deduplication pipeline for financial data begins with URL filtering, in which the raw data is initially processed. This crucial step ensures the inclusion of only pertinent URLs, enhancing the dataset’s quality by excluding irrelevant or unsuitable sources. Once the URLs are streamlined, the Text extraction phase commences, whereby contents of documents from the selected URLs are meticulously extracted, filtering out images while maintaining the large dataset scale. Following this, the language identification phase excluded non-English documents by categorizing them based on the language of their tokens. Subsequently, the pipeline further refines the data through Document-wise domain-based filtering, narrowing down to 100 billion tokens pertinent to the financial domain by excluding 55B-token non-financial documents. Recognizing the importance of data privacy and relevance, the pipeline incorporates removing sensitive information, which is done using a classifier built using FinBERT Araci (2019). Line-wise corrections enhance accuracy and filter out 5B tokens of sensitive information. An extensive Fuzzy deduplication process reduces the data to 38 billion tokens. This is followed by an Exact deduplication method, which trims another 13 billion tokens. Finally, the text cleaning process identifies and excludes 5B improper tokens, including all sensitive information. Ultimately, the pipeline crafts a streamlined financial dataset, culminating in a concise 20B-token financial dataset. The pipeline is illustrated in figure 3.
@@ Appendix D Downstream Dataset details
Figure D.1: Illustration of the FinVQA Dataset to understand model’s proficiency in Chart Understanding Tasks. The figure demonstrates the model’s ability to interpret stock price movements and identify peak values from a candlestick chart.
@@
Retrieval Data extraction Answer Generation From the following PDF file, extract all the relevant information that might help in answering the question: PDF: URL Extracted Text: text…….. Question: ask question? You are a financial expert specializing in the nuanced analysis of financial statements and a wide array of data-driven financial tasks. And think step by step for each prompt. Sometimes, you must extract relevant information to proceed with the problem. Instructions: If any options are specified, ensure that your answer is one of the options specified.
new Do not explain why you think the answer is correct.
new Answer the question by formulating your response using predefined mathematical functions. For addition, use Add(a, b), which represents a + b. Use Subtract(a, b) to denote a - b for subtraction. Construct your answer by combining these functions appropriately to reflect the required calculations.
Context: text + table + image
Retrieval: Retrieved relevant information Question: Task based question? Answer:
@@
You are a financial expert specializing in the nuanced analysis of financial statements and a wide array of data-driven financial tasks. And think step by step for each prompt. Sometimes, you must extract relevant information to proceed with the problem. Instructions: new If any options are specified, ensure that your answer is one of the options specified.
new Do not explain why you think the answer is correct.
@@
Figure D.4: FinTra Prompting method Example
@@
Hallucinations index Examples of No Hallucinations Prompt: What is the definition of ’Low/No Documentation Loan’ in the financial domain? Wikipedia: A No-Doc or Low-doc loan (abbr: No/Low Documentation Loan) refers to loans that do not require borrowers to provide income documentation to lenders or require little documentation. FinTral-RL-T&R: A low/no documentation loan allows a potential borrower to apply for a mortgage while providing little or no information regarding their employment, income, or assets. Regulation of these loans has evolved significantly since 2008, but they remain an option for some borrowers in nontraditional financial situations. Annotator Rating: A ChatGPT: A low/no documentation loan is a type of mortgage or loan where the borrower is not required to provide extensive financial documentation, such as income and asset verification, making it easier to qualify but often resulting in higher interest rates. Annotator Rating: A GPT-4: A Low/No Documentation Loan is a type of loan that requires minimal or no verification of the borrower’s income, assets, or employment, often used by individuals who may not have traditional income sources or documentation. Annotator Rating: A Examples of LLMs Hallucinations Prompt: What is the definition of ’Open Market’ in the financial domain? Wikipedia: An open market is an economic system without trade barriers to free market activities. In an open market, buyers and sellers can do business freely without common barriers, such as unfair licensing agreements, arbitrary taxes, unionization, subsidies, and other regulations affecting regular market operations. FinTral-RL-T&R: An open market is an economic system with little to no barriers to free-market activity. An open market is characterized by the absence of tariffs, taxes, licensing requirements, subsidies, unionization, and any other regulations or practices that interfere with free-market activity. Open markets may have competitive barriers to entry, but never any regulatory barriers to entry. Annotator Rating: A ChatGPT: In finance, the Open Market refers to a marketplace where securities are traded directly between buyers and sellers without the intervention of a third party, such as an exchange. Annotator Rating: C GPT-4: An "Open Market" in the financial domain refers to an economic system with few barriers to free market activity. It typically involves unrestricted buying and selling of goods and services, with prices determined by supply and demand, and minimal government intervention in terms of tariffs, quotas, or regulations. This concept is fundamental in the operation of capital markets, where securities are freely traded. Annotator Rating: A